# Recent Advances in Named Entity Recognition: A Comprehensive Survey and Comparative Study

### v2

@@ 1 Introduction
Named Entity Recognition (NER) is a subfield of computer science and Natural Language Processing (NLP) that focuses on identifying and classifying entities in unstructured text into predefined categories, such as persons, geographical locations and organizations (Grishman and Sundheim, 1996a). Over time, NER has expanded its scope beyond proper names to include more complex concepts, particularly in specialized domains such as biomedicine. For example, in the biomedical field, NER techniques are used to identify entities such as genes, proteins, and diseases (Mesbah et al., 2018; Wu et al., 2024; Luo et al., 2023; Mehmood et al., 2023; Lee et al., 2004). As a result, NER has become a crucial component in various modern applications, including machine translation (Babych and Hartley, 2003), question-answering systems (Mollá et al., 2006), and information retrieval (Guo et al., 2009).
Historically, early NER systems relied on rule-based approaches that used hand-crafted rules, lexicons, and spelling features (Rau, 1991; Mikheev et al., 1999; Farmakiotou et al., 2000). These methods were popular for their simplicity and interpretability, despite lacking flexibility and scalability. The introduction of machine learning techniques marked a significant shift in the field, allowing for more adaptable and data-driven approaches (Chieu and Ng, 2003; Nadeau and Sekine, 2007; Konkol and Konopík, 2013; Shaalan, 2014a; Eltyeb and Salim, 2014). With the rise of neural networks, NER systems further improved, particularly with the adoption of deep learning methods, which enabled more sophisticated models capable of capturing complex patterns in text (Collobert, 2011; Zhao et al., 2016; Zhang and Yang, 2018). Most recently, Transformer-based architectures have set new standards in NER performance, leading to breakthroughs in the field (Labusch et al., 2019; Jeong and Kim, 2022; Vacareanu et al., 2024; Martin et al., 2022).
This progress is reflected in the substantial growth of NER research publications in the last three decades. As shown in Figure 1, the number of NER-related publications in the ACM Computing Surveys database has grown steadily. In the mid-1990s, NER publications were relatively few, reflecting the field’s early focus on rule-based systems. The introduction of statistical models such as hidden Markov models (Baum et al., 1970) and conditional random fields (Vapnik, 1964; Cortes and Vapnik, 1995) around the year 2000 brought a surge in academic interest. More recently, the period 2018-2024 saw an explosion in NER publications, driven by the adoption of Transformer-based models that dramatically improved the performance of NER systems.
@@
To track how the field is evolving and provide structured overviews of advances that are made, a number of surveys have been proposed that document the progression from early rule-based methods to modern machine learning approaches (Nadeau and Sekine, 2007; Shaalan, 2014b). For example, Goulart et al. (2011) specifically reviewed advances in biomedical NER between 2007 and 2009, while Marrero et al. (2013) provided a broader theoretical and practical review of NER techniques. However, many of these early surveys predate the widespread adoption of deep learning and Transformer-based models. More recent surveys have shifted their focus to these modern approaches. For instance, several reviews have concentrated on deep learning and Transformer-based methods (Pakhale, 2023; Jehangir et al., 2023; Li et al., 2020a; Nasar et al., 2021; Dai et al., 2023; Li et al., 2022). Despite their depth, these surveys often overlook recent advances, such as the integration of Large Language Models (LLMs) and graph-based techniques. Research using graph-based methods, such as the work by Wang et al. (2022a), focuses on specific challenges such as nested entities but does not address flat-named entities. Similarly, biomedical NER surveys, such as Wang et al. (2023a), mention the use of pre-trained language models, but lack detailed analysis of their application in this domain.
A detailed review of NER in historical documentation by Ehrmann et al. (2023) examines the distinct challenges and strategies relevant to this field. This research provides significant insights into NER applied to historical texts, which often present unique issues such as language evolution, non-standardization, and inconsistent spelling. However, while it addresses these specific challenges in historical documents, the review does not comprehensively cover broader advancements in NER technologies, such as LLMs or graph-based approaches, nor does it extensively discuss strategies for managing limited annotations. Moreover, several recent works that use LLMs for NER are not yet covered in existing surveys. For instance, Wang et al. (2023b) improve few-shot NER in new domains by incorporating type-specific features, while Ashok and Lipton (2023) uses entity type definitions to enhance few-shot learning.
@@ 2 Task definition
In a formal context, consider $T$ a sequence of $N$ tokens represented by $T=(t_{1},t_{2},\dots,t_{N})$. NER entails generating a set of tuples $(I_{s},I_{e},\ell)$, where $s$ and $e$ are integers confined to the interval $[1,N]$. Here, $I_{s}$ and $I_{e}$ denote respectively the start and end indices of mentions of a named entity, and $\ell$ indicates the category, from among a set of predefined categories, to which the entity belongs. For instance, in the sentence “Barack Obama was born in Honolulu.”, NER would identify “Barack Obama” as a person’s name and “Honolulu” as a location, as shown in Figure 2.
@@ 3.1 Nested named entities
Nested named entities are entities contained within other entities. For instance, in the sentence "Barack Obama was born in Honolulu, Hawaii." the term "Honolulu, Hawaii" is a location entity nested within another (Honolulu being the capital of the island state of Hawaii). Identifying such nested entities is crucial in specialized fields such as biomedical text mining, where entities frequently overlap and are intricately embedded within one another. The management of nested entities generally involves hierarchical models or multi-level tagging methods (Alex et al., 2007; Shibuya and Hovy, 2020; Wang et al., 2022b).
@@ 4 Applications of NER
In this section we present several illustrative examples of applications for NER.
@@
  • Information extraction: NER can be used to extract structured data from unstructured text, for example in retrieving the names of persons, organizations, or medical concepts. Studies such as (Nadeau and Sekine, 2007; Weston et al., 2019) have explored different ways of improving NER systems. To this end, Etzioni et al. (2005a) examined the use of unsupervised learning approaches, while Weston et al. (2019) investigated the use of deep learning to increase the accuracy and efficiency of entity recognition. These studies are examples of the progress currently being made in increasing the precision of information extraction.
  • Information retrieval: NER can significantly improve both traditional and conversational information retrieval (IR) by accurately identifying and classifying entities in user queries and responses. This improves the precision of search results and allows systems to better understand user intent. In IR, improvements come from precise identification of pertinent entities in both search queries and the resulting data (Banerjee et al., 2019; Cheng et al., 2021). Research by Cowie and Lehnert (1996) and Etzioni et al. (2005a) confirms that NER improves IR systems. In conversational systems, NER enables accurate understanding of context and user intent. For instance, when a user inquires about the weather at a specific location, NER extracts the location entity, allowing the conversational assistant to provide an appropriate response. Studies such as (Jayarao et al., 2018; Cheng et al., 2019; Park et al., 2023) highlight the effectiveness of NER in improving the performance of virtual assistants-.
@@
  • Question answering: NER can play a role in answering questions that require pinpointing particular entities (Mollá et al., 2006). For instance, in a question like "When did Steve Jobs die?", NER can identify "Steve Jobs" as a person and extract the death date. Works by Mollá et al. (2006) and Zhang et al. (2016) demonstrate the value of NER in improving question answering systems.
@@ 5.1 Knowledge-based methods
Knowledge-based methods have been a foundational part of NER, particularly in the early stages of the field. These methods, which originated from linguistic principles, rely on predefined rules and lexical resources to identify named entities. For example, Borkowski (1966) presented an algorithm that uses rule-based lists and lexical markers, such as capitalization patterns, to identify company names. This rule-driven approach allows systems to detect entities based on consistent patterns, such as prefixes like "Mr." or "Ms." that signal the presence of a named entity. Another notable example is the CasEN transducer cascade (Maurel et al., 2011), designed specifically to recognize French-named entities.
These methods, which do not require annotated data, are primarily based on rules and gazetteers. Gazetteers serve as essential resources, providing a collection of domain-specific entities that improve recognition performance. Figure 4 shows a typical architecture used in knowledge-based NER systems. Architectures generally involve three key components: (1) a set of rules for identifying named entities, (2) optional gazetteers for additional context, and (3) an extraction engine that applies the rules and gazetteers to the input text. Despite the effort required to build these systems “manually”, they offer robust performance in well-defined domains (Zhang and Elhadad, 2013a; Btoush et al., 2016) and when using well-adapted gazetteers (Etzioni et al., 2005b; Sekine and Nobata, 2004).
@@
One of the key applications of knowledge-based methods is in biomedical NER, where systems such as ProMiner (Hanisch et al., 2005) have been developed to identify gene and protein names in texts. These systems use synonym gazetteers and specialized detection procedures to address challenges such as synonym ambiguity and case sensitivity. Similarly, Quimbaya et al. (2016) applied a gazetteer-based technique to electronic health records, obtaining an improvement in recall with minimal impact on precision (refer to Section 8.2).
Other studies have applied knowledge-based methods in a variety of domains. For example, a rule-based approach was used to extract dietary recommendations from unstructured text, demonstrating the flexibility and domain-specific applicability of such methods (Eftimov et al., 2017). Hybrid methods that combine rule-based approaches with machine learning have been shown to improve the accuracy of entity recognition in complex texts such as biomedical literature (Zhang and Elhadad, 2013b). A comprehensive approach that combined rules and deep learning techniques for pseudonymizing clinical documents (Tannier et al., 2024) serves as a good illustration of the importance and the challenges of deidentification in clinical data warehouses.
An interesting extension to knowledge-based approaches involves the use of factored sequence labeling to extract methodology components from AI research articles. Ghosh et al. (2023) propose a data-driven factored sequence labeling approach that leverages both ontology-based and data-driven techniques. The ontology-based technique uses predefined categories from knowledge bases such as PaperswithCode, while the data-driven variant employs clustering on sentence embeddings to dynamically identify emerging methodologies. This combined approach allows for more precise extraction of scientific concepts, particularly in dynamic fields where terminology evolves rapidly. The factorized method captures dependencies between methodology names and their contexts, thereby overcoming the limitations of traditional rule-based approaches, especially for newly introduced concepts.
@@ 5.2 Feature-engineering-based methods
As NER evolved, the need to automate the extraction of named entities led to the development of methods based on feature engineering. The idea behind these approaches was to reduce manual rule-setting by focusing on identifying within the text key features that could be used for NER tasks. Feature-engineering-based methods can be broadly categorized into unsupervised, supervised, and semi-supervised learning methods, each with its own advantages and limitations.
@@ 5.2.1 Unsupervised learning methods
Shinyama and Sekine (2004a) demonstrated this principle by utilizing word distributions to identify named entities, particularly in news articles where named entities frequently co-occur. This method capitalizes on the tendency of named entities to appear together in multiple documents, distinguishing them from common nouns. Bonnefoy et al. (2011), calculated a semantic proximity score by comparing word distributions in documents linked to an entity and the type of entity. Similarly, Shinyama and Sekine (2004b) observed that named entities often appear synchronously in different news reports, whereas common nouns do not. Nadeau et al. (2006) introduced an unsupervised system to build gazetteers and to resolve the ambiguity of named entities. Inspired by previous works such as (Etzioni et al., 2005a), their approach involves amalgamating extracted gazetteers with publicly accessible gazetteers, and the resulting performance has been commendable. Unsupervised learning methods nonetheless face several limitations. First, the absence of supervision makes it difficult to assess the accuracy of the extracted named entities, where labeled data could potentially provide a clearer assessment of accuracy. Second, although word distribution methods capture named entities well, they may fail in disambiguating entities with similar surface forms but with different meanings. Finally, the complexity of syntagmatic structures means that unsupervised methods might overlook nuanced semantic differences between entities, leading to less precise groupings or associations.
@@ 5.2.2 Semi-supervised learning methods
One approach that is frequently adopted in semi-supervised NER involves co-training (Kozareva et al., 2005), where multiple classifiers are trained on different data views, and these classifier then serve to iteratively label unlabeled instances in order to build a more robust model. An alternative approach involves self-training (Gao et al., 2021), where the initial model is trained on the labeled data and then used to output the labels for the unlabeled data. The most confident predictions are added to the labeled data and the process is iterated.
Semi-supervised learning methods offer numerous advantages, such as reducing reliance on annotated datasets and improving outcomes in low-resource contexts. However, they also present challenges, among which we may mention the risk of propagating errors from the initial labeled data, and the requirement that handle noisy or inaccurate predictions from the unlabeled data are handled meticulously. Despite these challenges, when used judiciously, semi-supervised learning approaches can significantly improve NER performance, and they can have clear benefits in the real world, where labeled data may be scarce or costly to acquire.
@@ 5.2.3 Supervised learning methods
Several prominent models have been developed for NER tasks. These include the Hidden Markov Model (HMM) (Baum et al., 1970), which uses probabilistic methods to predict the sequence of labels (Zhou and Su, 2002; Morwal et al., 2012; Bikel et al., 1999a; Zhao, 2004). The Maximum Entropy model (MaxEnt) (Berger et al., 1996) focuses on finding the most likely label for a given word or phrase based on its context (Borthwick, 1999; Curran and Clark, 2003; Chieu and Ng, 2003; Lin et al., 2004). Support Vector Machines (SVM) are used for classification tasks, separating named entities from non-named entities with a clear margin (Isozaki and Kazawa, 2002; Makino et al., 2002; Ekbal and Bandyopadhyay, 2010; Ju et al., 2011). Finally, Conditional Random Fields (CRF) (Vapnik, 1964; Cortes and Vapnik, 1995) consider the context of the entire sentence to make predictions about named entities (McCallum and Li, 2003; Settles, 2004a; Nongmeikapam et al., 2011).
@@
This equation posits that given a sequence of tokens $T$, the probability of identifying the sequence of named entities $N$ conditional on $T$ is the same as the probability of identifying the sequence of hidden states $S$ conditional on observations $O$. Bikel et al. (1999b) introduced IdentiFinder, an early HMM-based system that effectively learns to identify and categorize names, dates, times, and numerical quantities. The methodology was developed further in subsequent studies such as (Zhou and Su, 2002; Morwal et al., 2012; Bikel et al., 1999a; Zhao, 2004).
@@
MaxEnt models correspond to an advanced statistical method often used in NLP tasks, including NER. The core principle of these models is to determine a probability distribution over potential outcomes by maximizing entropy, subject to a predefined set of empirical constraints. The resulting distribution, characterized by having the highest entropy allowable under the observed constraints, is uniquely determined and coincides with the maximum likelihood distribution. It can be expressed as follows:
@@
CRFs are probabilistic models (Lafferty et al., 2001) widely used for sequence labeling tasks, such as sentence annotation. CRFs take account of the interdependence of neighboring elements, which makes these models exceptionally effective for NER. By capturing sequential dependencies among tokens within a sequence, CRFs adeptly encapsulate the complex contextual relationships characteristic of named entities. Through the integration of both local and global information, CRFs facilitate the prediction of entity labels by considering adjacent token labels, thereby significantly reducing labeling ambiguity. A CRF model is expressed as follows:
@@
The NER problem can be formulated as a CRF, where the observations are processed strings, and the labels correspond to the possible named entities. The best sequence of named entities will thus correspond to some sequence of tokens, and finding this best sequence of named entities is equivalent to finding the best sequence of labels, i.e., argmax $P(Y=y|X=x)$. Shishtla et al. (2008) implemented a system that extracts information from research articles using CRF. They investigated regularization problems using the Gaussian model and focused on the efficient use of feature space with CRF. Settles (Settles, 2004b) presented a framework to recognize biomedical entities using CRF with a variety of features. They demonstrated that a CRF with only simple orthographic features could achieve good performances.
@@
Thus, while SVMs might not in themselves model sequence dependencies as well as CRFs do, careful feature engineering and a strategic use of contextual information can nevertheless enable them to perform NER tasks competently.
@@
NER systems will often use separate supervised approaches in combination. Srihari (2000) described a hybrid strategy drawing upon MaxEnt, HMM, and custom grammatical rules. Rule-based tagging is used for predictable patterns such as time and monetary expressions, while statistical models such as HMM handle variable entities such as names, locations, and organizations. MaxEnt is used in conjunction with features enriched by external gazetteers to improve tagging accuracy.
@@
Hybrid methods have proven particularly valuable in low-resource languages, where they can effectively leverage both linguistic rules and machine learning models to improve NER accuracy. For instance, a hybrid NER system for Punjabi, proposed by Bajwa and Kaur (2015), integrates rule-based methods with HMM. Developed without an existing dataset, the system involved manual tagging to create training and testing data under linguistic supervision. Two versions of NER were introduced: one using only HMM, and another combining HMM with handcrafted rules. Similarly, to address the lack of existing resources for Arabic, Shaalan and Oudah (2014) proposed a hybrid NER system combining rule-based methods with machine learning approaches. This system recognizes 11 types of entities, including Person, Location, and Organization, by integrating decision trees (Quinlan, 1986), SVM, and logistic regression classifiers (Cox, 1958).

### v3

@@ 1 Introduction
Named Entity Recognition (NER) is a subfield of computer science and Natural Language Processing (NLP) that focuses on identifying and classifying entities in unstructured text into predefined categories, such as persons, geographical locations, and organizations (Grishman and Sundheim, 1996). Over time, NER has expanded its scope beyond proper names to include more complex concepts (Mehmood et al., 2023), particularly in specialized domains such as biomedicine. For example, in the biomedical field, NER techniques are employed to identify entities such as genes, proteins, and diseases (Mesbah et al., 2018; Luo et al., 2023). Consequently, NER has become a crucial component in various modern applications, including machine translation (Babych and Hartley, 2003), question-answering systems (QA) (Mollá et al., 2006), and information retrieval (IR) (Guo et al., 2009).
Historically, early NER systems relied on rule-based approaches with hand-crafted rules, lexicons, and spelling features (Rau, 1991; Mikheev et al., 1999; Farmakiotou et al., 2000). These methods, while simple and interpretable, lacked flexibility and scalability. The introduction of machine learning techniques marked a significant shift in the field, allowing more adaptable and data-driven approaches (Chieu and Ng, 2003; Nadeau and Sekine, 2007; Konkol and Konopík, 2013; Shaalan, 2014; Eltyeb and Salim, 2014). With the rise of neural networks, NER systems further improved, particularly with the adoption of deep learning methods, which enabled more sophisticated models capable of capturing complex patterns in text (Collobert, 2011; Zhao et al., 2016; Zhang and Yang, 2018). Most recently, Transformer-based architectures have set new standards in NER performance, leading to breakthroughs in the field (Labusch et al., 2019; Jeong and Kim, 2022; Vacareanu et al., 2024; Shi and Kimura, 2024).
This progress is reflected in the substantial growth of NER research publications in the last three decades. As shown in Figure 1, the number of NER-related publications in the ACM Computing Surveys database has grown steadily. In the mid-1990s, NER publications were relatively few, reflecting the field’s early focus on rule-based systems. The introduction of statistical models such as hidden Markov models (Baum et al., 1970) and conditional random fields (Lafferty et al., 2001) around the year 2000 brought a surge in academic interest. More recently, the period 2018-2024 saw an explosion in NER publications, driven by the adoption of Transformer-based models that improved the performance of NER systems.
@@
To track the evolution of NER and provide structured overviews of recent advances, several surveys have been published documenting the progression from rule-based methods to modern machine learning approaches (Nadeau and Sekine, 2007; Shaalan, 2014; Goulart et al., 2011; Marrero et al., 2013). For example, Goulart et al. (2011) specifically reviewed advances in biomedical NER between 2007 and 2009, while Marrero et al. (2013) provided a broader theoretical and practical review of NER techniques. However, many of these early surveys predate the widespread adoption of deep learning and Transformer-based models. More recent surveys have shifted their focus to these modern approaches. For example, several reviews have concentrated on deep learning and Transformer-based methods (Pakhale, 2023; Jehangir et al., 2023; Li et al., 2020a; Nasar et al., 2021; Dai et al., 2023; Li et al., 2022). Despite their depth, these surveys often overlook recent advances, such as the integration of Large Language Models (LLMs) and graph-based techniques. Research using graph-based methods, such as the work by Wang et al. (2022a), focuses on specific challenges such as nested entities but does not address flat-named entities. Similarly, biomedical NER surveys, such as Wang et al. (2023a), mention the use of pre-trained language models, but lack detailed analysis of their application in this domain.
A detailed review of NER in historical documentation by Ehrmann et al. (2023) examines the distinct challenges and strategies relevant to this field. This research provides significant insights into NER applied to historical texts, which often present unique issues such as language evolution, non-standardization, and inconsistent spelling. However, while it addresses these specific challenges in historical documents, the review does not comprehensively cover broader advancements in NER technologies, such as LLMs or graph-based approaches, nor does it extensively discuss strategies to manage limited annotations. Moreover, several recent works that use LLMs for NER have not yet been covered in existing surveys. For example, Wang et al. (2023b) improve few-shot NER in new domains by incorporating type-specific features, while Ashok and Lipton (2023) uses entity type definitions to enhance few-shot learning.
@@ 2 Task definition
In a formal context, consider $T$ a sequence of $N$ tokens represented by $T=(t_{1},t_{2},\dots,t_{N})$. NER entails generating a set of tuples $(I_{s},I_{e},\ell)$, where $s$ and $e$ are integers confined to the interval $[1,N]$. Here, $I_{s}$ and $I_{e}$ denote respectively the start and end indices of mentions of a named entity, and $\ell$ indicates the category, from among a set of predefined categories, to which the entity belongs. For example, in the sentence “Barack Obama was born in Honolulu.”, NER would identify “Barack Obama” as a person’s name and “Honolulu” as a location, as shown in Figure 2.
@@ 3.1 Nested named entities
Nested named entities are entities contained within other entities. For example, in the sentence "Barack Obama was born in Honolulu, Hawaii." the term "Honolulu, Hawaii" is a location entity nested within another (Honolulu being the capital of the island state of Hawaii). Identifying such nested entities is crucial in specialized fields such as biomedical text mining, where entities frequently overlap and are intricately embedded within one another. The management of nested entities generally involves hierarchical models or multi-level tagging methods (Alex et al., 2007; Shibuya and Hovy, 2020; Wang et al., 2022a; Shen et al., 2021).
@@ 4 Applications of NER
In this section, we present several illustrative examples of applications for NER.
@@
  • Information extraction: NER can be used to extract structured data from unstructured text, for example in retrieving the names of persons, organizations, or medical concepts. Studies such as (Nadeau and Sekine, 2007; Weston et al., 2019) have explored different ways of improving NER systems. To this end, Etzioni et al. (2005) examined the use of unsupervised learning approaches, while Weston et al. (2019) investigated the use of deep learning to increase the accuracy and efficiency of entity recognition. These studies are examples of the progress currently being made in increasing the precision of information extraction.
  • Information retrieval: NER can significantly improve both traditional and conversational information retrieval (IR) by accurately identifying and classifying entities in user queries and responses. This improves the precision of search results and allows systems to better understand user intent. In IR, improvements come from precise identification of pertinent entities in both search queries and the resulting data (Banerjee et al., 2019; Cheng et al., 2021). Research by Cowie and Lehnert (1996) and Etzioni et al. (2005) confirms that NER improves IR systems. In conversational systems, NER enables accurate understanding of context and user intent. For example, when a user inquires about the weather at a specific location, NER extracts the location entity, allowing the conversational assistant to provide an appropriate response. Studies such as (Jayarao et al., 2018; Cheng et al., 2019; Park et al., 2023) highlight the effectiveness of NER in improving the performance of virtual assistants-.
@@
  • Question answering: NER can play a role in answering questions that require pinpointing particular entities (Mollá et al., 2006). For example, in a question like "When did Steve Jobs die?", NER can identify "Steve Jobs" as a person and extract the death date. Works by Mollá et al. (2006) and Zhang et al. (2016) demonstrate the value of NER in improving question answering systems.
@@ 5.1 Knowledge-based methods
Knowledge-based methods have been a foundational part of NER, particularly in the early stages of the field. These methods, which originate from linguistic principles, rely on predefined rules and lexical resources to identify named entities. For example, Borkowski (1966) presented an algorithm that uses rule-based lists and lexical markers, such as capitalization patterns, to identify company names. This rule-driven approach allows systems to detect entities based on consistent patterns, such as prefixes like "Mr." or "Ms." that signal the presence of a named entity. Another notable example is the CasEN transducer cascade (Maurel et al., 2011), designed specifically to recognize French-named entities.
These methods, which do not require annotated data, are primarily based on rules and gazetteers. Gazetteers serve as essential resources, providing a collection of domain-specific entities that improve recognition performance. Figure 4 shows a typical architecture used in knowledge-based NER systems. Architectures generally involve three key components: (1) a set of rules for identifying named entities, (2) optional gazetteers for additional context, and (3) an extraction engine that applies the rules and gazetteers to the input text. Despite the effort required to build these systems “manually”, they offer robust performance in well-defined domains (Btoush et al., 2016; Eftimov et al., 2017) with well-adapted gazetteers (Hanisch et al., 2005; Sekine and Nobata, 2004).
@@
One of the key applications of knowledge-based methods is in biomedical NER, where systems such as ProMiner (Hanisch et al., 2005) have been developed to identify gene and protein names in texts. These systems use synonym gazetteers and specialized detection procedures to address challenges such as synonym ambiguity and case sensitivity. Similarly, Quimbaya et al. (2016) applied a gazetteer-based technique to electronic health records, obtaining an improvement in recall with minimal impact on precision (see Section 8.2).
Other studies have applied knowledge-based methods in a variety of domains. For example, a rule-based approach was used to extract dietary recommendations from unstructured text, demonstrating the flexibility and domain-specific applicability of such methods (Eftimov et al., 2017). Hybrid methods that combine rule-based approaches with machine learning have been shown to improve the accuracy of entity recognition in complex texts such as biomedical literature (Zhang and Elhadad, 2013). A comprehensive approach that combined rules and deep learning techniques for pseudonymizing clinical documents (Tannier et al., 2024) serves as a good illustration of the importance and challenges of deidentification in clinical data warehouses.
An interesting extension to knowledge-based approaches involves the use of factored sequence labeling to extract methodology components from AI research articles. Ghosh et al. (2023) propose a data-driven factored sequence labeling approach that leverages both ontology-based and data-driven techniques. The ontology-based technique uses predefined categories from knowledge bases such as PaperswithCode, while the data-driven variant employs clustering on sentence embeddings to dynamically identify emerging methodologies. This combined approach allows for more precise extraction of scientific concepts, particularly in dynamic fields where terminology evolves rapidly. The factorized method captures the dependencies between methodology names and their contexts, thus overcoming the limitations of traditional rule-based approaches, especially for newly introduced concepts.
@@ 5.2 Feature-engineering-based methods
As NER evolved, the need to automate the extraction of named entities led to the development of methods based on feature engineering. The idea behind these approaches was to reduce manual rule-setting by focusing on identifying key features that could be used for NER tasks within the text. Feature-engineering-based methods can be broadly categorized into unsupervised, supervised, and semi-supervised learning methods, each with its own advantages and limitations.
@@ 5.2.1 Unsupervised learning methods
Shinyama and Sekine (2004) demonstrated this principle by utilizing word distributions to identify named entities, particularly in news articles where named entities frequently co-occur. This method capitalizes on the tendency of named entities to appear together in multiple documents, distinguishing them from common nouns. Bonnefoy et al. (2011), calculated a semantic proximity score by comparing word distributions in documents linked to an entity and the type of entity. Nadeau et al. (2006) introduced an unsupervised system to build gazetteers and to resolve the ambiguity of named entities. Inspired by previous works such as (Etzioni et al., 2005), their approach involves amalgamating extracted gazetteers with publicly accessible gazetteers, and the resulting performance has been commendable. However, unsupervised learning methods face several limitations. First, the absence of supervision makes it difficult to assess the accuracy of the extracted named entities, where labeled data could potentially provide a clearer assessment of accuracy. Second, although word distribution methods capture named entities well, they may fail in disambiguating entities with similar surface forms but with different meanings. Finally, the complexity of syntagmatic structures means that unsupervised methods might overlook nuanced semantic differences between entities, leading to less precise groupings or associations.
@@ 5.2.2 Semi-supervised learning methods
One approach that is frequently adopted in semi-supervised NER involves co-training (Kozareva et al., 2005), where multiple classifiers are trained on different data views, and these classifier then serve to iteratively label unlabeled instances in order to build a more robust model. An alternative approach involves self-training (Gao et al., 2021), where the initial model is trained on the labeled data and then used to generate the labels for the unlabeled data. The most confident predictions are added to the labeled data and the process is iterated.
Semi-supervised learning methods offer numerous advantages, such as reducing the dependence on annotated datasets and improving outcomes in low-resource contexts. However, they also present challenges, among which we may mention the risk of propagating errors from the initial labeled data, and the requirement that handling noisy or inaccurate predictions from the unlabeled data is handled meticulously. Despite these challenges, when used judiciously, semi-supervised learning approaches can significantly improve NER performance, and they can have clear benefits in the real world, where labeled data may be scarce or costly to acquire.
@@ 5.2.3 Supervised learning methods
Several prominent models have been developed for NER tasks. These include the Hidden Markov Model (HMM) (Baum et al., 1970), which uses probabilistic methods to predict the sequence of labels (Zhou and Su, 2002; Morwal et al., 2012; Bikel et al., 1999; Zhao, 2004). The Maximum Entropy model (MaxEnt) (Berger et al., 1996) focuses on finding the most likely label for a given word or phrase based on its context (Borthwick, 1999; Curran and Clark, 2003; Chieu and Ng, 2003; Lin et al., 2004; Borthwick et al., 1998). Support Vector Machines (SVM) are used for classification tasks, separating named entities from non-named entities with a clear margin (Cortes and Vapnik, 1995; Makino et al., 2002; Ekbal and Bandyopadhyay, 2010; Ju et al., 2011). Finally, Conditional Random Fields (CRF) (Lafferty et al., 2001) consider the context of the entire sentence to make predictions about named entities (McCallum and Li, 2003; Settles, 2004; Nongmeikapam et al., 2011; Shishtla et al., 2008).
@@
This equation posits that given a sequence of tokens $T$, the probability of identifying the sequence of named entities $N$ conditional on $T$ is the same as the probability of identifying the sequence of hidden states $S$ conditional on observations $O$. Bikel et al. (1999) introduced IdentiFinder, an early HMM-based system that effectively learns to identify and categorize names, dates, times, and numerical quantities. The methodology was further developed in subsequent studies, such as (Zhou and Su, 2002; Morwal et al., 2012; Zhao, 2004).
@@
MaxEnt models correspond to an advanced statistical method that is often used in NLP tasks, including NER. The core principle of these models is to determine a probability distribution over potential outcomes by maximizing entropy, subject to a predefined set of empirical constraints. The resulting distribution, characterized by having the highest entropy allowable under the observed constraints, is uniquely determined and coincides with the maximum likelihood distribution. It can be expressed as follows:
@@
CRFs are probabilistic models widely used for sequence labeling tasks, such as sentence annotation. CRFs take account of the interdependence of neighboring elements, which makes these models exceptionally effective for NER. By capturing sequential dependencies among tokens within a sequence, CRFs adeptly encapsulate the complex contextual relationships characteristic of named entities. Through the integration of both local and global information, CRFs facilitate the prediction of entity labels by considering adjacent token labels, thereby significantly reducing labeling ambiguity. A CRF model is expressed as follows:
@@
The NER problem can be formulated as a CRF, where the observations are processed strings, and the labels correspond to the possible named entities. The best sequence of named entities will thus correspond to some sequence of tokens, and finding this best sequence of named entities is equivalent to finding the best sequence of labels, i.e., argmax $P(Y=y|X=x)$. Shishtla et al. (2008) implemented a system that extracts information from research articles using CRF. They investigated regularization problems using the Gaussian model and focused on the efficient use of feature space with CRF. Settles (Settles, 2004) presented a framework to recognize biomedical entities using CRF with a variety of features. They demonstrated that a CRF with only simple orthographic features could achieve good performances.
@@
Thus, while SVMs might not model sequence dependencies in themselves as well as CRFs do, careful feature engineering and a strategic use of contextual information can nevertheless enable them to perform NER tasks competently.
@@
NER systems will often use separate supervised approaches in combination. Srihari (2000) described a hybrid strategy based on MaxEnt, HMM, and custom grammatical rules. Rule-based tagging is used for predictable patterns such as time and monetary expressions, while statistical models such as HMM handle variable entities such as names, locations, and organizations. MaxEnt is used in conjunction with features enriched by external gazetteers to improve tagging accuracy.
@@
Hybrid methods have proven particularly valuable in low-resource languages, where they can effectively leverage both linguistic rules and machine learning models to improve NER accuracy. For example, a hybrid NER system for Punjabi, proposed by Bajwa and Kaur (2015), integrates rule-based methods with HMM. Developed without an existing dataset, the system involved manual tagging to create training and testing data under linguistic supervision. Two versions of NER were introduced: one using only HMM, and another combining HMM with hand-crafted rules. Similarly, to address the lack of existing resources for Arabic, Shaalan and Oudah (2014) proposed a hybrid NER system combining rule-based methods with machine learning approaches. This system recognizes 11 types of entity, including Person, Location, and Organization, by integrating decision trees (Quinlan, 1986), SVM, and logistic regression classifiers (Cox, 1958).